you want to borrow $7600 to purchase a 2013 Camry what will be your monthly payment if the loan is for four years at 6.45% interest
Answer #1
PVOrdinary Annuity = C*[(1-(1+i/(f*100))^(-n*f))/(i/(f*100))]
C = Cash flow per period
i = interest rate
n = number of payments I f = frequency of payment
7600= Cash Flow*((1-(1+ 6.45/1200)^(-4*12))/(6.45/1200))
Cash Flow = 180.06
Using Calculator: press buttons "2ND"+"FV" then assign
PV =-7600
I/Y =6.45/12
N =4*12
FV = 0
CPT PMT
Using Excel
=PMT(rate,nper,pv,fv,type)
=PMT(6.45/(12*100),12*4,,7600,)